EBD* Blues Workshops and Dance

Saturday, November 6

The Democracy Center, Harvard Square 45 Mount Auburn St., Cambridge

Workshops 1:30 – 6 p.m. Dance 8:00 – 11 p.m.

Sunday practice session 2:30 – 4:30 p.m.

Register on-line for discounted admission

    www.beantownbluesdance.com

*EBD stands for “Everything But Dance,” but there will most definitely be dancing!

EBD* Blues Workshops and Dance

Saturday, November 6

The Democracy Center, Harvard Square 45 Mount Auburn St., Cambridge

Workshops 1:30 – 6 p.m. Dance 8:00 – 11 p.m.

Sunday practice session 2:30 – 4:30 p.m.

Register on-line for discounted admission

    www.beantownbluesdance.com

*EBD stands for “Everything But Dance,” but there will most definitely be dancing!

EBD* Blues Workshops and Dance

Saturday, November 6

The Democracy Center, Harvard Square 45 Mount Auburn St., Cambridge

Workshops 1:30 – 6 p.m. Dance 8:00 – 11 p.m.

Sunday practice session 2:30 – 4:30 p.m.

Register on-line for discounted admission

    www.beantownbluesdance.com

*EBD stands for “Everything But Dance,” but there will most definitely be dancing!

EBD* Blues Workshops and Dance

Saturday, November 6

The Democracy Center, Harvard Square 45 Mount Auburn St., Cambridge

Workshops 1:30 – 6 p.m. Dance 8:00 – 11 p.m.

Sunday practice session 2:30 – 4:30 p.m.

Register on-line for discounted admission

    www.beantownbluesdance.com

*EBD stands for “Everything But Dance,” but there will most definitely be dancing!